Thinfilms morning report 6/18/03

james haydon jhaydon at snf.stanford.edu
Wed Jun 18 09:06:48 PDT 2003


Thinfilms morning report  6/18/03



Balzers: ok
Gryphon: ok
Implanter: ok Problem with Arc current control
Innotec: ok
Metalica: ok. Down for regen later this morning should be back mid to
late afternoon
STS Dep: May have problems with low frequency oxide index of refraction
check your process
wbSOLVENT: ok
wbgen-CTB: ok
wbgen-LHP: ok
wbgen-RHP: ok
wbMETAL: ok
wbNITRIDE: ok
wbNONMETAL: ok
wbDIF: ok
wbSILICIDE: ok
wbGAAS-LHP: ok
wbGAAS-RHP: ok
wbMISCRES: ok
wbOPTRES: ok
CPD:  ok New piece of equipment
DI WATER: ok
NITROGEN: ok








More information about the thinfilms mailing list